OpenStudy (anonymous):

If someone could please work this out with me, i really want to understand this, I'm not good at math at all. thank you (: Solve: 4(x + 1) = 4x + 4 No Solution 0 All Real Numbers 1

OpenStudy (mrnood):

expand the bracket on the left hand side - and rewrite the formula in your next reply

OpenStudy (anonymous):

what do you mean by expanding?

OpenStudy (lucaz):

expand 4(x+1)= 4*x + 4*1= 4x+4 the equation now is 4x+4=4x+4 this means that for any value of x the equation is true, because you have the same thing on both sides, so the solution is all R

OpenStudy (anonymous):

/.\ i guess i kinda of understand. thank you @lucaz

OpenStudy (lucaz):

yw